OpenStudy (anonymous):

A survey about the student government program at a school finds that of 245 students: 218 students like the program 27 students think the program is unnecessary If a circle graph were made from this data, what would the measure of the central angle be for the group that likes the program? Round your answer to the nearest whole number.

OpenStudy (anonymous):

would you show me how you got it i really don't get this

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

i don't think so but i am trying.....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you have the answer so that i can check whether i am right???

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

ok....so should i still show you my working????

OpenStudy (anonymous):

how would you get it how did you get the answer

OpenStudy (anonymous):

i would like to see it i will look it up and try to get it inmy book right now

OpenStudy (anonymous):

ok....the total no. of angles in a circle is 360 degrees. therefore you can represent the angle we are finding by theta. \[\frac{ \theta }{ 360^{0} } \times 245 = 218\]

OpenStudy (anonymous):

so what i wrote in a mathematical expression simply means that the angle occupied by those who like the program over the total no. of angles of the total no. of students = the number of people who like the program...

OpenStudy (anonymous):

yes thank you for helping me i am still looking it up in my book i will let you now when i am done thank you so much for helping me again i am still trying to get this

OpenStudy (anonymous):

know

OpenStudy (anonymous):

okay.....goodluck! you are welcome :)